A solution is made by dissolving 0.652 mol0.652 mol of nonelectrolyte solute in 861 g861 g of benzene. Calculate the freezing point, ?f,Tf, and boiling point, ?b,Tb, of the solution.

Thank you in advance
Answer #1

molality of solution = moles of solute / mass of solution in kg = 0.652 / 0.861= 0.757 m

Depression in freezing point = Kf * molality = 1.86 * 0.757 = 1.408 0C

Freezing point = actual freezing point - Depression in freezing point = 5.5 - 1.408 = 4.59 0C

elevation in boiling point = Kb * molality = 2.53 * 0.757 = 1.92 0C

Boiling point = actual boiling point + elevation = 80.1 + 1.92 = 82.02 0C